OpenStudy (anonymous):

how can i find the number of a percent for example 20% of 45

OpenStudy (anonymous):

9

OpenStudy (anonymous):

is 0.9 the same as 9

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

No, wait, 0.09 is the same

OpenStudy (anonymous):

i wanna know how you can find it

OpenStudy (anonymous):

"percent" pretty much means "over 100," and 9/100=.09

OpenStudy (anonymous):

You find it wit hproportions

OpenStudy (anonymous):

20/100=x/45

OpenStudy (anonymous):

cross multiply 20*45=100x

OpenStudy (anonymous):

then solve for x 900=100x 9=x

OpenStudy (anonymous):

so 9 is 20% of 45

OpenStudy (anonymous):

thanks